# MolTC: Towards Molecular Relational Modeling In Language Models

### v5

@@ B.3 Modules
In our experiments, the two graph encoder are instantiated by the five-layer GINE Hu et al. (2019). We conduct 2 million molecules from the ZINC15 Sterling and Irwin (2015) dataset to pretrain them by contrastive learning following Liu et al. (2023). Similarly, two projector are initialized with the encoder-only transformer, Sci-BERT, which is pretrained on scientific publications Beltagy et al. (2019), while its cross-attention modules are randomly initialized. More detailed pretraining process of our Q-Formors follows the training process in Liu et al. (2023), such as there are 8 query tokens in Q-Formers ($N_{q}=8$). Note that for LLM-based baselines, we fine-tune the backbone LLMs on task-specific datasets for fair comparison. Their prediction is considered accurate only if the outputs include words or numbers that correctly depict the interaction in question, without presenting any that describe alternative interactions.
@@ B.5 Training Strategy
We employ the AdamW optimizer Loshchilov and Hutter (2017) with a weight decay set at 0.05. Our learning rate strategy utilizes a combination of linear warm-up and cosine decay, optimizing the training process by initially increasing the learning rate to promote faster convergence, and then gradually decreasing it according to a cosine curve to fine-tune the model parameters. LoRA is implemented using the Open Delta library Ding et al. (2022), and the PEFT library Mangrulkar et al. (2022). LoRA’s rank $r$ is set to 16, while LoRA is applied to Galactica’s modules of [q_proj,v_proj,out_proj,fc1, fc2] following Liu et al. (2023). This configuration yields a LoRA adapter with 12M parameters which constitutes merely 0.94% of the parametersin the $\text{Galactica}_{\text{1.3B}}$.
@@ Appendix C Future Work
In this paper, we introduce a novel unified framework, leveraging LLM technology to predict molecular interactive properties. The future development directions of this project are twofold. First, there is an emphasis on expanding its application scope, for instance, applying it to downstream tasks such as few-shot learning. Second, we aim to enhance its capabilities by incorporating technologies like graph explainability Fang et al. (2023a, b), graph sampling Wang et al. (2022, 2023); Fang et al. (2024), and spatio-temporal modeling Xia et al. (2023); Wu et al. (2023), making it more comprehensive or enabling it to process multiple inputs simultaneously, instead of just two.

### v6

@@ B.3 Modules
In our experiments, the two graph encoder are instantiated by the five-layer GINE Hu et al. (2019). We conduct 2 million molecules from the ZINC15 Sterling and Irwin (2015) dataset to pretrain them by contrastive learning following Liu et al. (2023a). Similarly, two projector are initialized with the encoder-only transformer, Sci-BERT, which is pretrained on scientific publications Beltagy et al. (2019), while its cross-attention modules are randomly initialized. More detailed pretraining process of our Q-Formors follows the training process in Liu et al. (2023a), such as there are 8 query tokens in Q-Formers ($N_{q}=8$). Note that for LLM-based baselines, we fine-tune the backbone LLMs on task-specific datasets for fair comparison. Their prediction is considered accurate only if the outputs include words or numbers that correctly depict the interaction in question, without presenting any that describe alternative interactions.
@@ B.5 Training Strategy
We employ the AdamW optimizer Loshchilov and Hutter (2017) with a weight decay set at 0.05. Our learning rate strategy utilizes a combination of linear warm-up and cosine decay, optimizing the training process by initially increasing the learning rate to promote faster convergence, and then gradually decreasing it according to a cosine curve to fine-tune the model parameters. LoRA is implemented using the Open Delta library Ding et al. (2022), and the PEFT library Mangrulkar et al. (2022). LoRA’s rank $r$ is set to 16, while LoRA is applied to Galactica’s modules of [q_proj, v_proj, out_proj, fc1, fc2] following Liu et al. (2023a). This configuration yields a LoRA adapter with 12M parameters which constitutes merely 0.94% of the parametersin the $\text{Galactica}_{\text{1.3B}}$.
@@ Appendix C Future Work
In this paper, we introduce a novel unified framework, leveraging LLM technology to predict molecular interactive properties. The future development directions of this project are twofold. First, we aim to adopt advanced graph encoding techniques to enhance the comprehension of molecular graphs Gao et al. (2023); Wang et al. (2023a); Yu et al. (2023b, 2024); Gao et al. (2024). Secondly, in light of the extensive input information introduced in multi-molecular input scenarios, we plan to employ techniques such as graph interpretability Ying et al. (2019); Luo et al. (2020); Zheng et al. (2023); Fang et al. (2024) and sparsification Wang et al. (2023b, c) to eliminate information redundancy and resist out-of-distribution Fang et al. (2023a, b).